OpenStudy (anonymous):

Conjecture: The sum of the first 30 odd numbers is _? I know the answer is 900 but can someone please explain to me the steps to find the answer?

OpenStudy (anonymous):

use the first term as 1 and the common difference as +2 and one can use AP formulas for summation

OpenStudy (anonymous):

Can you go a little more in depth with that explanation please?

OpenStudy (anonymous):

here the number of terms is 30 Sum upto n terms =(30/2)((2*1) +(30-1)*2)=30/2*60=900